# Evidence Contextualization and Counterfactual Attribution for Conversational QA over Heterogeneous Data with RAG Systems

### v2

@@ Abstract.
Retrieval Augmented Generation (RAG) works as a backbone for interacting with an enterprise’s own data via Conversational Question Answering (ConvQA). In a RAG system, a retriever fetches passages from a collection in response to a question, which are then included in the prompt of a large language model (LLM) for generating a natural language (NL) answer. However, several RAG systems today suffer from two shortcomings: (i) retrieved passages usually contain their raw text and lack appropriate document context, negatively impacting both retrieval and answering quality; and (ii) attribution strategies that explain answer generation usually rely only on similarity between the answer and the retrieved passages, thereby only generating plausible but not causal explanations. In this work, we demonstrate RAGonite, a RAG system that remedies the above concerns by: (i) contextualizing evidence with source metadata and surrounding text; and (ii) computing counterfactual attribution, a causal explanation approach where the contribution of an evidence to an answer is determined by the similarity of the original response to the answer obtained by removing that evidence. To evaluate our proposals, we release a new benchmark ConfQuestions, with $300$ hand-created conversational questions, each in English and German, coupled with ground truth URLs, completed questions, and answers from $215$ public Confluence pages, that are typical of enterprise wiki spaces with heterogeneous elements. Experiments with RAGonite on ConfQuestions show the viability of our ideas: contextualization improves RAG performance, and counterfactual attribution is effective at explaining RAG answers.
@@ 1. Introduction
Motivation. Talking to your data is a major research theme today, where a user interacts with local knowledge repositories to satisfy their information needs. Conversational question answering (ConvQA) (Reddy et al., 2019; Christmann et al., 2019) is a natural choice for such interactions, where a user starts with a self-sufficient (intent-explicit) question, and follows that up with more ad hoc, conversational questions that leave parts of the context unspecified (intent-implicit). Hand in hand, the emergence of powerful large language models (LLMs) has led to retrieval augmented generation (RAG) (Lewis et al., 2020) as the backbone for designing QA systems over one’s own data. In RAG pipelines, given a question, a retriever fetches relevant evidence from local data, and this is passed on to an LLM for a concise and fluent answer to the user’s question. Enterprise knowledge repositories usually consist of documents with heterogeneous elements, i.e. they often contain interleaved tables, lists and passages (henceforth, we use the unifying term evidence). Typical examples of such documents with mixed structured and unstructured elements are meeting notes, test reports, or product descriptions. The setting for this demo is ConvQA with RAG over such heterogeneous document collections (Pan et al., 2022; Yu, 2022).
@@
The benchmark and all other artifacts for this work are public at https://gitlab.cc-asp.fraunhofer.de/iis-nlp-public/ragonite-public and https://github.com/Fraunhofer-IIS/RAGonite.
@@ 2. System overview
Backend. An overview of the RAGonite pipeline is in Fig. 1. RAGonite’s backend is split into a functional core, which handles retrieval and answer generation, and a stateful layer that persists chats into a SQLite database and provides a REST API to the frontend using FastAPI. The dependencies in the functional core include the vector database (ChromaDB), and prompt template (Jinja) and LLM libraries (gpt-4o). We use ChromaDB as our vector database for storing the contextualized evidences, and also use ChromaDB’s in-built retrieval functions. The multilingual BGE embedding model [^1] https://huggingface.co/BAAI/bge-m3 was used to embed the documents, that was found to work slightly better than text-embedding-3-small [^2] https://openai.com/index/new-embedding-models-and-api-updates/ from OpenAI. While ChromaDB was used for its extensibility, we also explored variants like Weaviate and Milvus. We used top-k hybrid search with the BGE re-ranker [^3] https://huggingface.co/BAAI/bge-reranker-v2-m3 being used to merge $k$ dense retrieval and lexical retrieval results ($k=10$ in our experiments). We use gpt-4o as our LLM of choice inside the question completion and answer generation modules for efficiency and quality, but we also support other OpenAI, Mistral, and Llama models. All prompts can be seen inside the demo interface for transparency. A single GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores) was used for all our experiments. All code is in Python.
@@ 3. Demo walkthrough
We use a screenshot of the RAGonite main page in Fig. 2, on which we overlay numbered blue boxes for a demo walkthrough. First, a user must select a domain (0) on which to use RAGonite (our focus here is on enterprise wikis exemplified by Confluence, but RAGonite also runs on other domains like soccer, automobiles, and fictional universes). Then the user uses the question input box (1) to begin a conversation. Suppose we are at turn two (2): when RAGonite receives a conversational question as input, it uses its question completion module to rephrase the question into an intent-explicit form using an LLM (3). While the question in the first conversation turn is usually self-contained, follow-up questions are completed using relevant information from previous questions and generated answers. Offline, the heterogeneous document collection is preprocessed via our evidence contextualization module (Sec. 4) into a pool of evidence where necessary document context is concatenated to the raw contents of the evidence (example in Fig. 3, where the page title and the preceding heading is prepended to a verbalized table record). A retriever then takes the intent-explicit question and searches over the evidence pool, to return top-k question-relevant evidences (4, 4a-c) using the hybrid retrieval score (5). These top-k evidences are inserted into the prompt of another LLM instance to generate the answer (6). Finally, the generated answer and the top-k evidences are fed into our counterfactual attribution module (Sec. 5). When a user clicks on “Explain” (6a), a command line window pops up to output the attribution distribution as explanation of how the answer was potentially constructed (Fig. 4). We log user feedback on the answer for future use (7). A user can see a trace through the pipeline using behind-the-scenes buttons (8): for example, retrieval results for lexical and dense search, LLM prompts, and more. Users can adjust retriever and generator configurations (9), see their past conversations (10), including deleted ones (11). Follow-up question suggestions are also provided (12) to help domain exploration. Notably, most open-source RAG demos only consist of the evidence retriever and answer generator modules. On average, RAGonite requires about one second to answer a question, and approximately two seconds for explaining the answer.
@@ 4. Evidence contextualization
Evidence contextualization in current RAG systems mainly involve steps like: (i) coreference resolution for resolving simple pronouns like (s)he, her, his, etc., and (ii) overlapping text chunking involving sliding windows over passages (Gao et al., 2023). These have some basic limitations like: (i) resolving pronouns is mainly limited to previous entity references, and is not enough for more subtle coreferences like “these configurations” or “the previous model”; (ii) sliding word/token windows do not generalize well to structured evidence like tables. We, however, adopt a simple but effective alternative of concatenating document context to each evidence at indexing time.
@@
We store each table in a verbalized form (Oguz et al., 2022), which converts structured evidence to a form more amenable to an LLM prompt while retaining scrutability by a human. In this mode, we traverse left to right in a table row, and linearize the content as “¡Column header 1¿ is ¡value 1¿, and ¡Column header 2¿ is ¡value 2¿, …”. We prepend this text with “Row ¡id¿ in Table ¡id¿” to preserve original row ordering information as well as table ordering relative to the page where the content is located. For instance, the second row in the table in Fig. 5 would be verbalized as “Row 2 in Table 1: Member is Alice, and Task is Similarity function, and …”. Along with the complete table, each linearized row is also stored as a unit evidence. This helps pinpoint answer provenance to individual rows in tables when possible, as well as resolve comparative questions that are easier when the whole table is in one piece in the prompt.
At this point, we have passages, lists, tables, and table records as individual evidence, ordered by their position in the page. We then concatenate the following items to each evidence: page title, previous heading, the evidence before, and the evidence after. Both the raw textual forms and their embeddings (text-embedding-3-small model from OpenAI) are indexed in our database, enabling lexical and dense search, respectively. Note that indexing whole pages and inserting full text of top-k documents into an LLM prompt cannot easily bypass evidence contextualization, due to arbitrary lengths of web documents: operating at evidence-level is a practical choice.
An example of contextualized evidence is in Fig. 3 (inside dashed box): the raw content of the evidence (the seventh row in the first table) starts with “Row 7 in Table 1…” But the preceding content starting at “OpenXT 9.0 …” consists of the page title, preceding heading, and preceding text. “OTA upgrade” at the end comes from adding succeeding text (the table footer here).
@@ 2 Output: Distribution $\boldsymbol{\mathcal{A}}$ for attributing answer $a$ to each evidence $e$
       compute $\boldsymbol{\mathcal{A}}\leftarrow softmax(s_{i})\forall e^{cl}_{i}$
@@
Idea. Counterfactual evidence through passage perturbation has been explored in the context of RAG (Sudhi et al., 2024) and more generally for LLM understanding (Xie et al., 2024), but not for answer attribution and not via evidence removal. The intuition here is that a definitive way of saying that how much an evidence $e_{i}$ was responsible for contributing to an answer is to remove the evidence from the prompt, and see how much the answer $a$ changed. If $a$ did not change much, then $e_{i}$ is unlikely to have played any major role, and vice versa.
@@
Counterfactual Attribution (CFA). This simple idea is the core of our proposed Algorithm 1. The only other confounding factor is the presence of redundant evidences in the prompt: the retriever may sometimes fetch two (or more) semantically equivalent evidences if they both satisfy the user’s intent. In the context of counterfactuals, it may be that when one of these redundant evidences are removed, the answer hardly changes. This may lead to the wrong conclusion that the first evidence was unimportant: it just happens that the second evidence makes up for it content-wise. Hence, it is important that we first cluster evidences $E$ (Line $3$) by content before we apply an iterative removal of each evidence cluster $e^{cl}_{i}$ (Lines $4$ and $5$). Moreover, to adjust for non-determinism in LLM generations, we use a Monte Carlo method: in this case it simply means that we repeat the answer generation $m$ times (Lines $6-7$), compute similarities between the answer and the counterfactual answer in each iteration $a^{cf}_{i,j}$ (Line $8$), and average these $m$ values for each cluster (Line $10$). This average similarity is the contribution of each evidence cluster to the original answer. These raw answer contribution scores $s_{i}$ for each evidence cluster are then normalized via softmax to derive the final attribution distribution $\mathcal{A}$ (Fig. 4). We used the popular DBSCAN algorithm for clustering evidences (Line 3) as it requires only two tunable parameter ($\epsilon$ and $minPts$), and does not require the apriori specification of the number of clusters. Cosine similarity between the text embeddings of the answer and the counterfactual answer (in both cases the completed question from the benchmark is prepended for providing the context in which the similarity is sought) was used in Line 8, and using JinaAI embeddings [^4] https://huggingface.co/jinaai/jina-embeddings-v3 to encode evidences which was found to work better than other embedding models in this context. While our algorithm produces fairly accurate results (Table 3), output distributions (example in Fig. 4) may sometimes appear to be somewhat degenerate to end users. For example, say among six clusters, the top-evidence cluster is attributed $20\%$ and the remaining five groups are all assigned about $16\%$. We thus introduce a temperature parameter $t$ to skew the distribution more towards the top evidence cluster. Answers and counterfactual answers were generated using the cheaper gpt4o-mini model (instead of gpt4o) due to budget constraints: every question entails the generation of up to ten counterfactual answers, which entails a substantially higher number of output tokens than generating the intent-explicit questions or the original answers.
@@
Parameter selection. Parameters $\epsilon$, $minPts$, and $t$ were set to $0.005,2$, and $0.05$, respectively, using domain and pipeline knowledge. Specifically, (i) the $\epsilon$ parameter in DBSCAN defines the maximum distance between two points for them to be considered neighbors and thus indirectly controls the number of clusters: setting it to be $0.005$ typically resulted in a maximum cluster size of about $2-3$ within $10$ evidences. This is optimal because we cannot attribute answers at a fine granularity if too many evidences get grouped together; (ii) we work with ten evidences, so $minPts$, the number of samples in a neighborhood for a point to be considered as a core point, inside the DBSCAN algorithm is set to two; and (iii) we observed that setting the the temperature to $0.05$ skews the distributions in apparently degenerate cases to $1-3$ evidences: this reflects the average number of gold evidences as per our benchmark.
@@
Differences from ContextCite. Very recently, we came to know about contemporary work on counterfactual attribution in RAG by Cohen-Wang et al. (Cohen-Wang et al., 2024) (referred to as ContextCite). Our ideas were independently developed and differ from ContextCite in four major aspects: (i) ContextCite only applies for models with open weights where answer generation likelihoods can be exactly computed, while our idea of estimating evidence attribution scores from similarities of answers and counterfactual answers makes it applicable to RAG systems with closed models as well; (ii) ContextCite does not deal with the problem of redundant information that we tackle via evidence clustering; and (iii) ContextCite operates via sentence-level ablations to create counterfactual scenarios: ubiquitous coreferences in sentences then becomes an unsolved concern. We preclude this problem by rather operating at evidence-level, where each evidence is contextualized as discussed earlier, and this makes most evidences relatively self-contained; and (iv) the evaluation of attribution is very different in ContextCite: while they rely on log-probability drops and rank correlations, our benchmark ConfQuestions allows for more precise evaluation using an accuracy metric.

### v3

@@ Abstract.
Retrieval Augmented Generation (RAG) works as a backbone for interacting with an enterprise’s own data via Conversational Question Answering (ConvQA). In a RAG system, a retriever fetches passages from a collection in response to a question, which are then included in the prompt of a large language model (LLM) for generating a natural language (NL) answer. However, several RAG systems today suffer from two shortcomings: (i) retrieved passages usually contain their raw text and lack appropriate document context, negatively impacting both retrieval and answering quality; and (ii) attribution strategies that explain answer generation typically rely only on similarity between the answer and the retrieved passages, thereby only generating plausible but not causal explanations. In this work, we demonstrate RAGonite, a RAG system that remedies the above concerns by: (i) contextualizing evidence with source metadata and surrounding text; and (ii) computing counterfactual attribution, a causal explanation approach where the contribution of an evidence to an answer is determined by the similarity of the original response to the answer obtained by removing that evidence. To evaluate our proposals, we release a new benchmark ConfQuestions: it has $300$ hand-created conversational questions, each in English and German, coupled with ground truth URLs, completed questions, and answers from $215$ public Confluence pages. These documents are typical of enterprise wiki spaces with heterogeneous elements. Experiments with RAGonite on ConfQuestions show the viability of our ideas: contextualization improves RAG performance, and counterfactual explanations outperform standard attribution.
@@ 1. Introduction
Motivation. “Talk to your data” is a major research theme today, where users interact with local knowledge repositories to satisfy their information needs. Conversational question answering (ConvQA) (Reddy et al., 2019; Christmann et al., 2019) is a natural choice for such interactions, where a user starts with a self-sufficient (intent-explicit) question, and follows that up with more ad hoc, conversational questions that leave parts of the context unspecified (intent-implicit). Hand in hand, the emergence of powerful large language models (LLMs) has led to retrieval augmented generation (RAG) (Lewis et al., 2020) as the backbone for designing QA systems over one’s own data. In RAG pipelines, given a question, a retriever fetches relevant evidence from local data, and this is passed on to an LLM for a concise and fluent answer to the user’s question. Enterprise knowledge repositories usually consist of documents with heterogeneous elements, i.e. they often contain interleaved tables, lists and passages (henceforth, we use the unifying term evidence). Typical examples of such documents with mixed structured and unstructured elements are meeting notes, test reports, or product descriptions. The setting for this demo is ConvQA with RAG over such heterogeneous document collections (Pan et al., 2022; Yu, 2022).
@@
The benchmark and all other artifacts for this work are public at https://github.com/Fraunhofer-IIS/RAGonite.
@@ 2. System overview
Backend. An overview of the RAGonite pipeline is in Fig. 1. RAGonite’s backend is split into a functional core, which handles retrieval and answer generation, and a stateful layer that persists chats into a SQLite database and provides a REST API to the frontend using FastAPI. The dependencies in the functional core include the vector database (ChromaDB), and prompt template (Jinja) and LLM libraries (gpt-4o and Llama-3.1-8B). We use ChromaDB as our vector database for storing the contextualized evidences, and also use ChromaDB’s in-built retrieval functions. The multilingual BGE embedding model [^1] https://huggingface.co/BAAI/bge-m3 was used to embed evidences, that was found to work slightly better than text-embedding-3-small [^2] https://openai.com/index/new-embedding-models-and-api-updates/ from OpenAI. While ChromaDB was used for its extensibility, we also explored variants like Weaviate and Milvus. We used top-k hybrid search with the BGE reranker [^3] https://huggingface.co/BAAI/bge-reranker-v2-m3 and reciprocal rank fusion (RRF) (Cormack et al., 2009) being used to merge $k$ dense retrieval and lexical retrieval results. We use GPT-4o as our LLM of choice inside the question completion and answer generation modules for efficiency and quality, but we also support a Llama model in our demo. All prompts can be seen inside the demo interface for transparency. A single GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores) was used for all our experiments. All code is in Python.
@@ 3. Demo walkthrough
We use a screenshot of the RAGonite main page in Fig. 2, on which we overlay numbered blue boxes, for a demo walkthrough. First, a user must select a domain (0) on which to use RAGonite (our focus here is on enterprise wikis exemplified by Confluence, but RAGonite also runs on other domains like soccer, automobiles, and fictional universes). Then the user uses the question input box (1) to begin a conversation. Suppose we are at turn two (2): when RAGonite receives a conversational question as input, it uses its question completion module to rephrase the question into an intent-explicit form using an LLM (3). While the question in the first conversation turn is usually self-contained, follow-up questions are completed using relevant information from previous questions and generated answers. Offline, the heterogeneous document collection is preprocessed via our evidence contextualization module (Sec. 4) into a pool of evidence where necessary document context is concatenated to the raw contents of the evidence (example in Fig. 3, where the page title and the preceding heading is prepended to a verbalized table record). A retriever then takes the intent-explicit question and searches over the evidence pool, to return top-k question-relevant evidences (4, 4a-c) using the hybrid retrieval score (5). These top-k evidences are inserted into the prompt of another LLM instance to generate the answer (6). Finally, the generated answer and the top-k evidences are fed into our counterfactual attribution module (Sec. 5). When a user clicks on “Explain” (6a), a command line window pops up to output the attribution distribution as an explanation of how the answer was potentially constructed (Fig. 4). We log user feedback on the answer for future use (7). A user can see a trace through the pipeline using behind-the-scenes buttons (8): for example, retrieval results for lexical and dense search, LLM prompts, and more. Users can adjust retriever and generator configurations (9), see their past conversations (10), including deleted ones (11). Follow-up question suggestions are also provided (12) to help domain exploration. Notably, most open-source RAG demos only consist of the evidence retriever and answer generator modules. On average, RAGonite requires about one second to answer a question, and approximately two seconds for explaining the answer.
@@ 4. Evidence contextualization
Idea. Evidence contextualization in current RAG systems mainly involve steps like: (i) coreference resolution for resolving simple pronouns like (s)he, her, his, etc., and (ii) overlapping text chunking involving sliding windows over passages (Gao et al., 2023). These have some basic limitations like: (i) resolving pronouns is mainly limited to previous entity references, and is not enough for more subtle coreferences like “these configurations” or “the previous model”; (ii) sliding word/token windows do not generalize well to structured evidence like tables. We, however, adopt a simple but effective alternative of concatenating document context to each evidence at indexing time.
@@
Preprocessing. We store each table in a verbalized form (Oguz et al., 2022), which converts structured evidence to a form more amenable to an LLM prompt while retaining scrutability by a human. In this mode, we traverse left to right in a table row, and linearize the content as “¡Column header 1¿ is ¡value 1¿, and ¡Column header 2¿ is ¡value 2¿, …”. We prepend this text with “Row ¡id¿ in Table ¡id¿” to preserve original row ordering information as well as table ordering relative to the page where the content is located. For instance, the second row in the table in Fig. 5 would be verbalized as “Row 2 in Table 1: Member is Alice, and Task is Similarity function, and …”. Along with the complete table, each linearized row is also stored as an individual evidence. This helps pinpoint answer provenance to individual rows in tables when possible, as well as resolve comparative questions that are easier when the whole table is in one piece in the prompt.
Contextualization. At this point, we have passages, lists, tables, and table records as individual evidences. We then concatenate the following items to each piece of evidence: page title, previous heading, the evidence before, and the evidence after. Both the raw textual forms and their embeddings are indexed in our database, enabling lexical and dense search, respectively. Note that indexing whole pages and inserting full text of top-k documents into an LLM prompt cannot easily bypass evidence contextualization, due to arbitrary lengths of web documents: operating at evidence-level is therefore a practical choice.
An example of contextualized evidence is in Fig. 3 (inside dashed box): the raw content of the evidence (the seventh row in the first table) starts with “Row 7 in Table 1…” But the previous content starting at “OpenXT 9.0 …” consists of the page title, preceding heading, and preceding text. “OTA upgrade” at the end comes from adding succeeding text (the table footer here).
@@ 2 Output: Distribution $\boldsymbol{\mathcal{A}}$ for attributing answer $a$ to each evidence $e$
       compute $\boldsymbol{\mathcal{A}}\leftarrow softmax(c_{i})\forall e^{cl}_{i}$
@@
Idea. Counterfactual explanations (Tran et al., 2021) through passage perturbation has been explored in the context of RAG (Sudhi et al., 2024) and more generally for LLM understanding (Xie et al., 2024), but not for answer attribution and not via evidence removal, as explore in this work. The intuition here is that a definitive way of saying how much an evidence $e_{i}$ contributed to an answer $a$ is to remove the evidence from the prompt, and see how much the original answer $a$ changed. If $a$ did not change much, then $e_{i}$ is unlikely to have played any major role, and vice versa.
@@
Counterfactual Attribution (CFA). This simple idea is the core of our proposed Algorithm 1. The only other confounding factor is the presence of redundant evidences in the prompt: the retriever may sometimes fetch two (or more) semantically equivalent evidences if they both satisfy the user’s intent. In the context of counterfactuals, it may be that when one of these redundant evidences are removed, the answer hardly changes. This may lead to the wrong conclusion that the first evidence was unimportant: it just happens that the second evidence makes up for it content-wise. Hence, it is important that we first cluster evidences $E$ (Line $3$) by content before we apply an iterative removal of each evidence cluster $e^{cl}_{i}$ (Lines $4$ and $5$). Each cluster’s removal and corresponding counterfactual answer generation is implemented in parallel inside the demo for efficiency, as evidences (clusters) can be processed independent of each other. Moreover, to adjust for non-determinism in LLM generations, we use a Monte Carlo method: in this case it simply means that we repeat the answer generation $m$ times (Lines $6-7$), compute similarities between the answer and the counterfactual answer in each iteration $a^{cf}_{i,j}$ (Line $8$), and average these $m$ values for each cluster. One minus this average similarity is the contribution of each evidence cluster to the original answer: the more similar the answers are, the less an evidence contributed (Line $10$). These raw answer contribution scores $c_{i}$ for each evidence cluster are then normalized via masked softmax (Line $11$) to derive the final attribution distribution $\mathcal{A}$ (Fig. 4). We used the popular DBSCAN algorithm for clustering evidences (Line 3) as it requires only two tunable parameters ($\epsilon$ and $minPts$), and does not require the specification of the number of clusters upfront (hard to predict in our case). Cosine similarity between the text embeddings of the answer and the counterfactual answer (in both cases the completed question from the benchmark is prepended for providing the context in which the similarity is sought) was used in Line 8. Using JinaAI embeddings [^4] https://huggingface.co/jinaai/jina-embeddings-v3 to encode evidences which was found to work better than other embedding models (BGE, OpenAI) in this context.
While our algorithm produces fairly accurate results (Table 3), output distributions (example in Fig. 4) may sometimes appear to be somewhat degenerate to end users. For example, say among six clusters, the top-evidence cluster is attributed $20\%$ and the remaining five groups are all assigned about $16\%$. We thus introduce a temperature parameter $t$ to introduce a higher skew of the distribution towards the top-scoring evidence cluster. Answers and counterfactual answers were generated using the cheaper GPT-4o-mini model (instead of GPT-4o) due to budget constraints: every question entails the generation of up to ten counterfactual answers, which entails a substantially higher number of output tokens than generating the intent-explicit questions or the original answers.
@@
Parameter selection. Parameters $\epsilon$, $minPts$, and $t$ were set to $0.005,2$, and $0.05$, respectively, using domain and pipeline knowledge. Specifically, (i) the $\epsilon$ parameter in DBSCAN defines the maximum distance between two points for them to be considered neighbors and thus indirectly controls the number of clusters: setting it to be $0.005$ typically resulted in a maximum cluster size of about $2-3$ within $10$ evidences. This is optimal because we cannot attribute answers at a fine granularity if too many evidences get grouped together; (ii) we work with just ten evidences, so $minPts$, the number of samples in a neighborhood for a point to be considered as a core point, inside the DBSCAN algorithm is set to two; and (iii) we observed that setting the the temperature to $0.05$ skews the distributions in apparently degenerate cases to $1-3$ evidences: this reflects the average number of gold evidences in ConfQuestions.
@@
Differences from ContextCite. Very recently, we came to know about contemporary work on counterfactual attribution in RAG by Cohen-Wang et al. (Cohen-Wang et al., 2024) (referred to as ContextCite). Our ideas were independently developed and differ from ContextCite in four major aspects: (i) ContextCite only applies for models with open weights where answer generation likelihoods can be exactly computed, while our idea of estimating evidence attribution scores from similarities of answers and counterfactual answers makes it applicable to RAG systems with closed models as well; (ii) ContextCite does not deal with the problem of redundant information that we tackle via evidence clustering; (iii) ContextCite operates via sentence-level ablations to create counterfactual scenarios: ubiquitous coreferences in sentences then becomes an unsolved concern. We preclude this problem by rather operating at evidence-level, where each evidence is contextualized as discussed earlier, and this makes most evidences relatively self-contained; and (iv) the evaluation of attribution is very different in ContextCite: while they rely on log-probability drops and rank correlations, our benchmark ConfQuestions allows for more precise evaluation using an accuracy metric.